OpenStudy (anonymous):

Sides of a quadrilateral are all positive integers . Three of them are 5,10,20 . How many possible values are there for the fourth side?

OpenStudy (anonymous):

5+10+20 = 35 in order to be a quadrilateral, the 4th side cannot be 35, otherwise you would have a straight line it also cannot be 0 otherwise you will have a triangle so it can only be integers from 1 to 34

OpenStudy (mayankdevnani):

The length of the fourth side can range from 20-10-5 = 5 to 20+10+5 = 35. These extreme values will give a "quadrilateral" that looks like a line segment. This range includes 31 possible values. If you insist it have non-zero area, the allowed values range from 6 to 34, a total of 29 possible values.